OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2)

OpenStudy (cwrw238):

expand the brackets 4x + 24 = 6x - 12 now subtract 4x from both sides, then add 12 to both sides and see what you get

OpenStudy (anonymous):

4x+24=6x-12 -2x=-36 x=18

OpenStudy (anonymous):

Thanks @cwrw238 I got the answer doing it the way you told me.

OpenStudy (cwrw238):

@pkjha3105 - it doesn't help much to give the answer to the problem. we are here to help people learn maths

OpenStudy (cwrw238):

yw

OpenStudy (jiteshmeghwal9):

4x+24=6x-12 4x-6x=-12-24 -2x=-36 x=12

OpenStudy (jiteshmeghwal9):

x=18*

OpenStudy (anonymous):

@cwrw238 one more if it doesnt bother how would i do one with a fraction like this: Solve: x + 3 = −x + 7

OpenStudy (anonymous):

**Solve: -1/2 x + 3 = −x + 7

OpenStudy (jiteshmeghwal9):

x+3=-x+7 x+x=7-3 2x=4 x=2

OpenStudy (anonymous):

Thank you!